# Decoding for Punctured Convolutional and Turbo Codes: A Deep Learning Solution for Protocols Compliance

### v1

@@ 3.1 Overall Architecture
The proposed Convolutional Neural Engine introduces a groundbreaking approach to decoding convolutional codes, particularly under punctured conditions, by effectively integrating domain-specific insights into a deep learning framework. Unlike conventional NN-based methods, which often treat punctured sequences as missing data, CNE incorporates an innovative Puncture Embedding mechanism that explicitly encodes puncturing patterns into the feature space. This approach enhances the decoder’s adaptability, enabling it to generalize across varying code rates specified by protocols and ensuring compatibility for practical applications. The architecture of the proposed CNE is illustrated in Figure 8.
@@ 3.1.1 Convolutional Decoding
At the core of the CNE architecture lies a sequence of operations that transform raw input data and puncturing patterns into actionable decoding outputs. The process begins with a projection of the input sequence, $\boldsymbol{x}\in\mathbb{R}^{K\times D_{\text{in}}}$, into a higher-dimensional embedding space:
$\boldsymbol{e}=\boldsymbol{W}_{\text{proj}}\boldsymbol{x}+\boldsymbol{b}_{\text{proj}}$ (7)
where $\boldsymbol{W}_{\text{proj}}\in\mathbb{R}^{D_{\text{embed}}\times D_{\text{in}}}$ and $\boldsymbol{b}_{\text{proj}}\in\mathbb{R}^{D_{\text{embed}}}$ are learnable parameters, $K$ represents the sequence length, and $D_{\text{in}}$ denotes the dimensionality of each input vector in the sequence. Specifically, for convolutional codes as defined in IEEE 802.11, $D_{\text{in}}=2$, corresponding to two input symbols per time step. Here, $\boldsymbol{x}$ represents the de-punctured data, with zeros inserted at the positions of punctured bits to restore the full sequence structure.
To incorporate puncturing information, a Puncture Embedding module maps the puncturing pattern $\boldsymbol{p}\in\mathbb{R}^{K\times D_{\text{in}}}$ into the same embedding space. This is achieved using:
$\boldsymbol{e}_{\text{p}}=\sigma\left(\boldsymbol{W}_{\text{punc}}\boldsymbol{p}+\boldsymbol{b}_{\text{punc}}\right)$ (8)
where $\boldsymbol{W}_{\text{punc}}\in\mathbb{R}^{D_{\text{embed}}\times D_{\text{in}}}$, $\boldsymbol{b}_{\text{punc}}\in\mathbb{R}^{D_{\text{embed}}}$, and $\sigma(\cdot)$ is the sigmoid activation function, which ensures that the puncture information acts as a gate, controlling the flow of puncturing effects.
The input embedding $\boldsymbol{e}$ and puncture embedding $\boldsymbol{e}_{\text{p}}$ are then combined through element-wise multiplication:
$\boldsymbol{e}_{\text{final}}=\boldsymbol{e}\odot\boldsymbol{e}_{\text{p}}$ (9)
@@
Following this, the combined embedding $\boldsymbol{e}_{\text{final}}$ undergoes batch normalization to stabilize training:
$\boldsymbol{e}_{\text{norm}}=\text{BN}(\boldsymbol{e}_{\text{final}})$ (10)
@@
$\boldsymbol{h}=\text{LSTM}(\boldsymbol{e}_{\text{norm}})$ (11)
where $\boldsymbol{h}\in\mathbb{R}^{K\times 2D_{\text{hidden}}}$ represents the encoded context of the entire sequence, leveraging information from both directions.
@@
$\boldsymbol{y}=\boldsymbol{W}_{\text{out}}\boldsymbol{h}+\boldsymbol{b}_{\text{out}}$ (12)
where $\boldsymbol{W}_{\text{out}}\in\mathbb{R}^{2D_{\text{hidden}}\times 1}$, and $\boldsymbol{y}\in\mathbb{R}^{K\times 1}$ represents the likelihoods of the estimated bits being 1.
@@
$\boldsymbol{y}=\boldsymbol{W}_{\text{out}}\cdot\text{LSTM}\Big{(}\text{BN}\Big{(}\big{(}\boldsymbol{W}_{\text{proj}}\boldsymbol{x}+\boldsymbol{b}_{\text{proj}}\big{)}\odot\sigma\big{(}\boldsymbol{W}_{\text{punc}}\boldsymbol{p}+\boldsymbol{b}_{\text{punc}}\big{)}\Big{)}\Big{)}+\boldsymbol{b}_{\text{out}}$ (13)
This unified formulation captures the seamless interaction among input embedding, puncture information, temporal modeling, and decoding. The proposed architecture’s explicit incorporation of puncturing patterns into the decoding process distinguishes it from traditional NN-based methods, demonstrating superior adaptability and performance in decoding convolutional codes.
@@ 3.1.2 Turbo Decoding
Before the iterative decoding process begins, the received sequence undergoes a rate-matching reversal procedure, which consists of de-puncturing and sub-block de-interleaving operations. During de-puncturing, zeros are inserted at the locations of punctured bits to restore the original codeword structure. These steps reconstruct the structure of the original transmitted codewords, ensuring compatibility with the decoder architecture. Additionally, this process generates the puncturing indicators $\boldsymbol{p}$, corresponding to the transmitted codeword and its parity sequences.
@@
In our designed architecture CNE 0 processes the systematic bit sequence $\boldsymbol{x}$ alongside the first parity sequence $\boldsymbol{z}$. CNE 1 processes the interleaved systematic bit sequence and the second parity sequence $\boldsymbol{z}^{\prime}$. These two CNEs are identical in architecture and parameterization, ensuring symmetry and simplifying training. Each CNE employs LSTM layers to extract temporal dependencies in the input sequence, followed by a fully connected layer to predict Log-Likelihood Ratios (LLRs).
@@
Step $\text{Step 1.}\quad\hat{\boldsymbol{x}}_{0}^{(t)}=f_{\text{CNE 0}}(\begin{bmatrix}\boldsymbol{x}+\hat{\boldsymbol{x}}_{1}^{\text{de-int},(t-1)},\boldsymbol{z}\end{bmatrix},\boldsymbol{p}_{0})$ (14)
Step $\text{Step 2.}\quad\hat{\boldsymbol{x}}_{0}^{\text{int},(t)}=\boldsymbol{\pi}(\hat{\boldsymbol{x}}_{0}^{(t)}-\hat{\boldsymbol{x}}_{1}^{\text{de-int},(t-1)})$
Step $\text{Step 3.}\quad\hat{\boldsymbol{x}}_{1}^{(t)}=f_{\text{CNE 1}}(\begin{bmatrix}\hat{\boldsymbol{x}}_{0}^{\text{int},(t)},\boldsymbol{z}^{\prime}\end{bmatrix},\boldsymbol{p}_{1})$
Step $\text{Step 4.}\quad\hat{\boldsymbol{x}}_{1}^{\text{de-int},(t)}=\boldsymbol{\pi}^{-1}(\hat{\boldsymbol{x}}_{1}^{(t)}-\hat{\boldsymbol{x}}_{0}^{\text{int},(t)})$
Here, $t$ denotes the iteration step, while $\pi$ and $\pi^{-1}$ represent the interleaving and de-interleaving operations, respectively. The vectors $\boldsymbol{p}_{0}$ and $\boldsymbol{p}_{1}$ indicate the puncturing information for $\begin{bmatrix}\boldsymbol{x},\boldsymbol{z}\end{bmatrix}$ and $\begin{bmatrix}\boldsymbol{x},\boldsymbol{z}^{\prime}\end{bmatrix}$, respectively.
After $N_{\text{iter}}$ iterations, the final estimate of the systematic bit likelihood is obtained by applying the de-interleaving operation to the output of CNE 1:
$\boldsymbol{y}=\boldsymbol{\pi}^{-1}\left(\hat{\boldsymbol{x}}_{1}^{(N_{\text{iter}})}\right)$ (15)
$\overline{\boldsymbol{\text{LLR}}}=\left\lvert\frac{\boldsymbol{\text{LLR}}-\mu}{\sqrt{\sigma^{2}+\epsilon}}\right\rvert\cdot\text{sign}(\boldsymbol{\text{LLR}})$ (16)
where $\mu$ and $\sigma^{2}$ denote the mean and variance of the LLR values, respectively, and $\epsilon=10^{-6}$ is a small constant added to ensure numerical stability by preventing division by zero. The function $\text{sign}(\cdot)$ preserves the sign of the LLRs during normalization.
@@ 3.2 Training Methodology
To optimize the proposed CNE-based convolutional and Turbo decoder for diverse code rates and puncturing patterns, a two-stage training strategy is employed. The process begins with pretraining on non-punctured codes at a fixed Signal-to-Noise Ratio (SNR) of 0 dB. This pretraining step establishes a strong foundational model capable of decoding convolutional and Turbo codes under idealized conditions, serving as the basis for subsequent fine-tuning.
@@
$\text{SNR}_{\text{train}}=\text{SNR}_{\text{offset}}+10\log_{10}(2R)$ (17)
where $\text{SNR}_{\text{offset}}$ serves as a baseline SNR value calibrated to align the Bit Error Rate (BER) across code rates. This parameter can be selected based on empirical results from conventional decoders, such as the Viterbi or BCJR algorithms, to ensure approximate BER alignment across code rates. This Balanced BER Training (BBT) prevents the loss function from being dominated by specific code rates, maintaining stability during optimization.
@@
The training process leverages the Adam optimizer [32] with an initial learning rate of $10^{-3}$ during pretraining and $10^{-4}$ during fine-tuning. The reduced learning rate in the fine-tuning phase helps preserve the knowledge acquired during pretraining and minimizes the risk of catastrophic forgetting. To further enhance optimization, a cosine decay scheduler [33] is employed to progressively decrease the learning rate to $10^{-6}$ by the end of training. Both pretraining and fine-tuning are carried out over 1000 epochs with a batch size of 128. The training and inference processes are performed on an NVIDIA RTX 4090 GPU with 24GB of memory, complemented by an Intel(R) Xeon(R) Platinum 8468V CPU and 512GB of DDR5 memory, providing the computational power necessary to handle the complexity and diversity of the dataset.
@@ 4 Experiments
In the experimental setup, the information bit length during training is 120, while during inference, it is varied across 120, 240, 480, and 960 to evaluate the model’s generalization to different input sizes. Training is performed in an AWGN channel, with inference conducted in both AWGN and Rayleigh fading channels to assess generalization across different channel conditions. Convolutional codes are assigned code rates of 1/2, 2/3, 3/4, and 5/6, in line with protocol specifications. For consistency, the Turbo codes are configured with code rates of 1/3, 1/2, 2/3, 3/4, and 5/6, as Turbo codes offer more flexibility in rate adaptation. The fine-tuning rates are 1/3, 1/2, 2/3, and 3/4, while the 5/6 code rate is used as an unseen rate only during inference to validate the model’s generalization ability. The traditional convolutional decoder uses the Viterbi algorithm with a traceback depth of 120, while the Turbo code decoder employs a reduced-complexity BCJR algorithm called max-log-MAP [34] with full traceback depth.
@@
The Rayleigh fading channel is modeled with 3 independent taps and a MIMO configuration with 4 transmit and 4 receive antennas. Channel State Information (CSI) in the Rayleigh fading scenario is obtained using least squares channel estimation [35]. For a received signal $\boldsymbol{y}$, the transmitted pilot matrix $\boldsymbol{P}$, and the channel matrix $\boldsymbol{H}$, the LS estimation is given by:
$\boldsymbol{\hat{H}}=\boldsymbol{y}\boldsymbol{P}^{\dagger}$ (19)
@@
where $\boldsymbol{P}^{\dagger}=(\boldsymbol{P}^{H}\boldsymbol{P})^{-1}\boldsymbol{P}^{H}$ is the Moore-Penrose pseudoinverse of $\boldsymbol{P}$, and $\boldsymbol{P}^{H}$ represents the Hermitian transpose of $\boldsymbol{P}$. The pilot matrix $\boldsymbol{P}\in\mathbb{C}^{N_{T}\times N_{T}}$ is generated according to the IEEE 802.11 protocol [18] to facilitate accurate channel estimation.
@@
For MIMO detection, the Minimum Mean Square Error (MMSE) algorithm is employed [36], which minimizes the mean squared error between the estimated transmitted signal $\hat{\boldsymbol{x}}$ and the true transmitted signal $\boldsymbol{x}$. The MMSE filter $\boldsymbol{W}\in\mathbb{C}^{N_{T}\times N_{R}}$ is computed as:
$\boldsymbol{W}=\left(\overline{\boldsymbol{H}}^{H}\overline{\boldsymbol{H}}+\lambda\boldsymbol{I}\right)^{-1}\overline{\boldsymbol{H}}^{H}$ (20)
@@
where $\lambda=10^{-6}$ is a regularization parameter ensuring numerical stability, and $\overline{\boldsymbol{H}}\in\mathbb{C}^{(N_{R}+N_{T})\times N_{T}}$ is the extended matrix incorporating the noise variance $\sigma^{2}$. The extended channel matrix is defined as:
$\overline{\boldsymbol{H}}=\begin{bmatrix}\boldsymbol{\hat{H}}\\ \sigma^{2}\boldsymbol{I}\end{bmatrix}$ (21)
where $\boldsymbol{I}\in\mathbb{C}^{N_{T}\times N_{T}}$ is the identity matrix.
Once the MMSE filter is computed, the transmitted signal $\boldsymbol{x}$ is estimated as:
$\hat{\boldsymbol{x}}=\boldsymbol{W}\boldsymbol{y}$ (22)
@@
This process yields the estimated transmitted signal $\hat{\boldsymbol{x}}$, which is subsequently passed through the demodulation and decoding stages to recover the original message.
@@
$\text{LLR}(x_{i})=\log\left(\frac{P(x_{i}=0\mid\hat{x})}{P(x_{i}=1\mid\hat{x})}\right)=\min_{\boldsymbol{x}_{0}}\left(\|\hat{x}-\boldsymbol{x}_{0}\|\right)-\min_{\boldsymbol{x}_{1}}\left(\|\hat{x}-\boldsymbol{x}_{1}\|\right)$ (23)
Here, $\boldsymbol{x}_{0}$ and $\boldsymbol{x}_{1}$ represent the possible transmitted symbols corresponding to hypotheses $x_{i}=0$ and $x_{i}=1$, respectively, while $\hat{x}$ denotes the detected signal.
@@
Following the demodulation process, the LLRs are passed to the proposed CNE neural network, which is characterized by several key hyperparameters. These include the number of LSTM layers $N_{\text{layer}}$, the embedding dimension $D_{\text{embed}}$, the hidden layer size $D_{\text{hidden}}$, and the number of iterations $N_{\text{iter}}$. Specifically, in the LSTM cell shown in Eq. 4, the weight matrices $\boldsymbol{W}_{f},\boldsymbol{W}_{i},\boldsymbol{W}_{o},\boldsymbol{W}_{c}\in\mathbb{R}^{D_{\text{embed}}\times(D_{\text{embed}}+D_{\text{hidden}})}$. A detailed summary of the parameters for the CNE neural network architecture and the simulation setup is provided in Table 1.

### v3

@@ 3.1 Overall Architecture
The proposed convolutional neural engine introduces a groundbreaking approach to decoding convolutional codes, particularly under punctured conditions, by effectively integrating domain-specific insights into a deep learning framework. Unlike conventional NN-based methods, which often treat punctured sequences as missing data, CNE incorporates an innovative puncturing-aware embedding mechanism that explicitly encodes puncturing patterns into the feature space. This approach enhances the decoder’s adaptability, enabling it to generalize across varying code rates specified by protocols and ensuring compatibility for practical applications. The architecture of the proposed CNE is illustrated in Figure 8.
@@ 3.1.1 Convolutional Decoding
Before decoding, the received log-likelihood ratios (LLRs) are rate-dematched by inserting zeros at the positions of punctured bits to restore the original codeword length $N$. The vectors $\mathbf{llr}_{z}\in\mathbb{R}^{K}$ and $\mathbf{llr}_{z^{\prime}}\in\mathbb{R}^{K}$, together with their corresponding puncturing indicators $\mathbf{p}_{z}\in\{0,1\}^{K}$ and $\mathbf{p}_{z^{\prime}}\in\{0,1\}^{K}$, are concatenated to form the matrix $\bm{L}_{\text{m}}\in\mathbb{R}^{K\times 2}$ and its associated puncturing pattern $\bm{P}\in\{0,1\}^{K\times 2}$. These matrices serve as the input to the convolutional neural engine.
At the core of the CNE architecture lies a sequence of operations that transform raw input data and puncturing patterns into actionable decoding outputs. The process begins with a projection of the input sequence $\bm{L}_{\text{m}}$ into a higher-dimensional embedding space:
$\bm{E}_{l}=\bm{W}_{\text{proj}}\bm{L_{\text{m}}}+\bm{b}_{\text{proj}}$ (7)
where $\bm{W}_{\text{proj}}\in\mathbb{R}^{D_{\text{embed}}\times D_{\text{in}}}$ and $\bm{b}_{\text{proj}}\in\mathbb{R}^{D_{\text{embed}}}$ are learnable parameters, $K$ represents the sequence length, and the input dimension $D_{\text{in}}=2$ corresponds to the LLRs of the two coded bits generated per time step in IEEE 802.11 convolutional codes.
A puncturing-aware embedding module integrates puncturing information by mapping the pattern $\bm{P}\in\{0,1\}^{K\times D_{\text{in}}}$, where each element specifies whether the corresponding position in $\bm{L}_{\text{m}}$ is punctured (0) or non-punctured (1), into the embedding space of $\bm{E}_{p}$. This is achieved using:
$\bm{E}_{p}=\phi\left(\bm{W}_{\text{punc}}\bm{P}+\bm{b}_{\text{punc}}\right)$ (8)
where $\bm{W}_{\text{punc}}\in\mathbb{R}^{D_{\text{embed}}\times D_{\text{in}}}$, $\bm{b}_{\text{punc}}\in\mathbb{R}^{D_{\text{embed}}}$, and $\phi(\cdot)$ is the sigmoid activation function, which ensures that the puncture information acts as a gate, controlling the flow of puncturing effects.
The input embedding $\bm{E}_{l}$ and puncturing-aware embedding $\bm{E}_{p}$ are then combined through element-wise multiplication:
$\bm{E}_{lp}=\bm{E}_{l}\odot\bm{E}_{p}$ (9)
@@
Following this, the combined embedding $\bm{E}_{lp}$ undergoes batch normalization (BN) to stabilize training:
$\bm{E}_{\text{norm}}=\text{BN}(\bm{E}_{lp})$ (10)
@@
$\bm{S}_{\text{out}}=\text{LSTM}(\bm{E}_{\text{norm}})$ (11)
where $\bm{S}_{\text{out}}\in\mathbb{R}^{K\times 2D_{\text{hidden}}}$ represents the encoded context of the entire sequence, leveraging information from both directions.
@@
$\mathbf{llr}_{u}=\bm{W}_{\text{out}}\bm{S}_{\text{out}}+\bm{b}_{\text{out}}$ (12)
where $\bm{W}_{\text{out}}\in\mathbb{R}^{2D_{\text{hidden}}\times 1}$, and $\mathbf{llr}_{u}\in\mathbb{R}^{K}$ represents the likelihoods of the estimated bits being 1.
@@
$\mathbf{llr}_{u}=\;\bm{W}_{\text{out}}\cdot\text{LSTM}\Big(\text{BN}\Big(\big(\bm{W}_{\text{proj}}\bm{L}_{\text{m}}+\bm{b}_{\text{proj}}\big)\odot\phi\big(\bm{W}_{\text{punc}}\bm{P}+\bm{b}_{\text{punc}}\big)\Big)\Big)+\bm{b}_{\text{out}}$ (13)
@@ 3.1.2 Turbo Decoding
Before the iterative decoding process begins, the received sequence undergoes a rate-matching reversal procedure, which consists of de-puncturing and sub-block de-interleaving operations. During de-puncturing, zeros are inserted at the locations of punctured bits to restore the original codeword structure. These processes generate the systematic sequence $\mathbf{llr}_{s}\in\mathbb{R}^{K}$, the first and second parity sequences $\mathbf{llr}_{z}\in\mathbb{R}^{K}$ and $\mathbf{llr}_{z^{\prime}}\in\mathbb{R}^{K}$, along with their corresponding puncturing indicators $\mathbf{p}_{s}\in\{0,1\}^{K}$, $\mathbf{p}_{z}\in\{0,1\}^{K}$, and $\mathbf{p}_{z^{\prime}}\in\{0,1\}^{K}$, which serve as inputs to the CNE-based Turbo code decoder.
@@
In our designed architecture CNE 0 processes the systematic sequence $\mathbf{llr}_{s}$ alongside the first parity sequence $\mathbf{llr}_{z}$. CNE 1 processes the interleaved systematic bit sequence and the second parity sequence $\mathbf{llr}_{z^{\prime}}$. The two CNEs share identical architectures and weights, ensuring symmetry and simplifying training.
@@
[…]
$\text{Step 2.}\quad\widehat{\mathbf{llr}}_{0}^{\text{int},(t)}=\bm{\pi}\left(\widehat{\mathbf{llr}}_{0}^{(t)}-\widehat{\mathbf{llr}}_{1}^{\text{de-int},(t-1)}\right)$
[…]
$\text{Step 4.}\quad\widehat{\mathbf{llr}}_{1}^{\text{de-int},(t)}=\bm{\pi}^{-1}\left(\widehat{\mathbf{llr}}_{1}^{(t)}-\widehat{\mathbf{llr}}_{0}^{\text{int},(t)}\right)$
Here, $t$ denotes the iteration step, while $\bm{\pi}$ and $\bm{\pi}^{-1}$ represent the interleaving and de-interleaving operations, respectively. After $N_{\text{iter}}$ iterations, the final estimate of the systematic bit likelihood is obtained by applying the de-interleaving operation to the output of CNE 1:
$\mathbf{llr}_{u}=\bm{\pi}^{-1}\left(\widehat{\mathbf{llr}}_{1}^{(N_{\text{iter}})}\right)$ (15)
3.1.3 Structural Differences Between CNE and DeepTurbo
Figure 10: The architecture of the DeepTurbo decoder.
Figure 10 illustrates the architecture of the DeepTurbo decoder [29] with a bidirectional gated recurrent unit (Bi-GRU) decoding core and two soft-in soft-out (SISO) modules, where $\mathbf{P}_{\text{in}}$ and $\mathbf{P}_{\text{out}}$ represent posterior information with a feature size of $F_{s}$. Although both the proposed CNE and DeepTurbo employ iterative decoding inspired by the BCJR algorithm, their architectures differ significantly in three key aspects, as outlined below:
  1. 1.
    Shared Weights: Unlike DeepTurbo, where the neural network weights for the two SISO modules are independent, our proposed CNE0 and CNE1 share the same set of weights. This design reduces the total number of parameters, thereby lowering storage requirements and improving computational efficiency, making the architecture more practical for resource-constrained environments.
  2. 2.
    Enhanced High-Dimensional Embedding: While DeepTurbo expands only the posterior information into a higher-dimensional space, our CNE-based Turbo decoder extends the systematic bits ($\mathbf{llr}_{s}$), as well as both parity sequences ($\mathbf{llr}_{z}$ and $\mathbf{llr}_{z^{\prime}}$), into a high-dimensional embedding space. This comprehensive embedding approach enhances the expressive power of the neural network, enabling it to better capture complex relationships within the codeword.
  3. 3.
    Puncturing-Aware Embedding Layer: A critical distinction of our work is the introduction of a novel puncturing-aware embedding layer, which DeepTurbo lacks. This layer explicitly incorporates puncturing patterns into the neural network’s latent space, enabling seamless generalization across various code rates. This feature ensures compatibility with practical communication protocols, such as IEEE 802.11 and 3GPP TS 36.212, where puncturing is prevalent.
The proposed CNE architecture overcomes DeepTurbo’s limitations in generalizing to diverse code lengths and rates, making it more suitable for modern communication requirements.
@@ 3.2 Training Methodology
To optimize the proposed CNE-based convolutional and Turbo decoder for diverse code rates and puncturing patterns, a two-stage training strategy is employed. The process begins with pre-training on non-punctured codes at a fixed SNR of 0 dB. This pre-training step establishes a strong foundational model capable of decoding convolutional and Turbo codes under idealized conditions, serving as the basis for subsequent fine-tuning.
@@
[…]
where $\text{SNR}_{\text{offset}}$ serves as a baseline SNR value calibrated to align the bit error rate across code rates. This parameter can be selected based on empirical results from conventional decoders, such as the Viterbi or BCJR algorithms, to ensure approximate BER alignment across code rates. This balanced BER training (BBT) prevents the loss function from being dominated by specific code rates, maintaining stability during optimization.
@@
The convolutional code decoder and Turbo code decoder are trained separately. Since both decoders are based on the CNE architecture, the convolutional code decoder can share the same CNE0 or CNE1 component with the Turbo code decoder for decoding, but with distinct weights to account for their different coding structures and requirements.
Table 1: Number of Samples in Training, Validation, and Testing Datasets
Pre-Training Fine-Tuning Validation Testing AWGN Rayleigh Convolutional Codes $1.6384\times 10^{7}$ $2288\times 10^{7} 1.1$$\times 10^{5} 1.1$$1\times 10^{6}$ $4.1\times 10^{6}$ Turbo Codes $1.6384\times 10^{7}$ $6384\times 10^{7} 1.1$$\times 10^{5} 1.1$$1\times 10^{6}$ $4.1\times 10^{6}$
The training, validation, and testing datasets are generated by creating original messages of length $K$, which are then encoded and modulated using BPSK. Noise is added, and the signals are passed through an AWGN channel, following Equation (5) to produce the received signals. In the pre-training phase, the dataset contains $128\times 128\times 1000=1.6384\times 10^{7}$ samples. For fine-tuning, datasets include convolutional codes with three code rates and Turbo codes with four code rates, each with $4.096\times 10^{6}$ samples. The validation dataset has $1\times 10^{4}$ samples per SNR level from 0 to 10 dB. The testing dataset covers two scenarios: AWGN channel with SNRs from 0 to 10 dB and $1\times 10^{5}$ samples per SNR, and Rayleigh fading channel with SNRs from 0 to 40 dB and $1\times 10^{5}$ samples per SNR. A summary of the datasets is presented in Table 1. To ensure the independence of training and inference datasets without overlap, different random seeds are used for data generation in these two phases, and a cross-check is performed to ensure that no training data is used for inference.
The training process leverages the Adam optimizer [35] with an initial learning rate of $10^{-3}$ during pre-training and $10^{-4}$ during fine-tuning. The reduced learning rate in the fine-tuning phase helps preserve the knowledge acquired during pre-training and minimizes the risk of catastrophic forgetting. To further enhance optimization, a cosine decay scheduler [36] is employed to progressively decrease the learning rate to $10^{-6}$ by the end of training. Both pre-training and fine-tuning were conducted for 1000 epochs, with each epoch comprising 128 minibatches of 128 samples each. The training and inference processes are performed on an NVIDIA RTX 4090 GPU with 24GB of memory, complemented by an Intel(R) Xeon(R) Platinum 8468V CPU and 512GB of DDR5 memory, providing the computational power necessary to handle the complexity and diversity of the dataset.
@@ 4 Experiments
In the experimental setup, the information block length during training is 120, while during inference, it is varied across 120, 240, 480, and 960 to evaluate the model’s generalization to different input sizes. Training is performed in an AWGN channel, with inference conducted in both AWGN and Rayleigh fading channels to assess generalization across different channel conditions. Convolutional codes are assigned code rates of 1/2, 2/3, 3/4, and 5/6, in line with protocol specifications. For consistency, the Turbo codes are configured with code rates of 1/3, 1/2, 2/3, 3/4, and 5/6, as Turbo codes offer more flexibility in rate adaptation. The fine-tuning rates are 1/3, 1/2, 2/3, and 3/4, while the 5/6 code rate is used as an unseen rate only during inference to validate the model’s generalization ability. For fine-tuning, $\text{SNR}_{\text{offset}}$ is set to 2.5 dB for convolutional decoders and 1.5 dB for Turbo decoders. The traditional convolutional decoder uses the Viterbi algorithm with a traceback depth of 120, while the Turbo code decoder employs a reduced-complexity BCJR algorithm called max-log-MAP [37] with full traceback depth. It should be noted that scaling the extrinsic information in the max-log-MAP algorithm can enhance decoding performance [38]. All simulations of the BCJR algorithm in this paper use the max-log-MAP without a scaling factor.
@@
The Rayleigh fading channel is modeled with 3 independent taps and a MIMO configuration with 4 transmit and 4 receive antennas. Channel state information (CSI) in the Rayleigh fading scenario is obtained using least squares channel estimation [39]. For a received signal $\bm{y}$, the transmitted pilot matrix $\bm{\Omega}$, and the channel matrix $\bm{H}$, the LS estimation is given by:
$\bm{\hat{H}}=\bm{y}\bm{\Omega}^{\dagger}$ (18)
@@
where $\bm{\Omega}^{\dagger}=(\bm{\Omega}^{H}\bm{\Omega})^{-1}\bm{\Omega}^{H}$ is the Moore-Penrose pseudoinverse of $\bm{\Omega}$, and $\bm{\Omega}^{H}$ represents the Hermitian transpose of $\bm{\Omega}$. The pilot matrix $\bm{\Omega}\in\mathbb{C}^{N_{T}\times N_{T}}$ is generated according to the IEEE 802.11 protocol [18] to facilitate accurate channel estimation.
@@
For MIMO detection, the minimum mean square error (MMSE) algorithm is employed [40], which minimizes the mean squared error between the estimated transmitted signal $\hat{\bm{x}}$ and the true transmitted signal $\bm{x}$. The MMSE filter $\bm{W}\in\mathbb{C}^{N_{T}\times N_{R}}$ is computed as:
$\bm{W}=\left(\overline{\bm{H}}^{H}\overline{\bm{H}}+\lambda\bm{I}\right)^{-1}\overline{\bm{H}}^{H}$ (19)
@@
where $\lambda=10^{-6}$ is a regularization parameter ensuring numerical stability, and $\overline{\bm{H}}\in\mathbb{C}^{(N_{R}+N_{T})\times N_{T}}$ is the extended matrix incorporating the noise variance $\sigma^{2}$. The extended channel matrix is defined as:
$\overline{\bm{H}}=\begin{bmatrix}\bm{\hat{H}}\\ \sigma^{2}\bm{I}\end{bmatrix}$ (20)
where $\bm{I}\in\mathbb{C}^{N_{T}\times N_{T}}$ is the identity matrix.
Once the MMSE filter is computed, the transmitted signal $\bm{x}$ is estimated as:
$\hat{\bm{x}}=\bm{W}\bm{y}$ (21)
@@
This process yields the estimated transmitted signal $\hat{\bm{x}}$, which is subsequently passed through the demodulation and decoding stages to recover the original message.
@@
$\text{LLR}(x_{i})=\log\left(\frac{P(x_{i}=0\mid\hat{x})}{P(x_{i}=1\mid\hat{x})}\right)=\min_{\bm{x}_{0}}\left(\|\hat{x}-\bm{x}_{0}\|\right)-\min_{\bm{x}_{1}}\left(\|\hat{x}-\bm{x}_{1}\|\right)$ (22)
Here, $\bm{x}_{0}$ and $\bm{x}_{1}$ represent the possible transmitted symbols corresponding to hypotheses $x_{i}=0$ and $x_{i}=1$, respectively, while $\hat{x}$ denotes the detected signal.
$\overline{\mathbf{llr}}=\left\lvert\frac{\mathbf{llr}-\mu}{\sqrt{\delta^{2}+\epsilon}}\right\rvert\odot\text{sign}(\bm{\mathbf{llr}})$ (23)
where $\mu$ and $\delta^{2}$ denote the mean and variance of the LLRs after demodulation but before de-rate matching, comprising both systematic and parity sequences, respectively, and $\epsilon=10^{-6}$ is a small constant added to ensure numerical stability by preventing division by zero. The function $\text{sign}(\cdot)$ preserves the sign of the LLRs during normalization.
@@
Following the normalization process, the LLRs are processed by the de-rate matching module and then passed to the proposed CNE neural network, which is characterized by several key hyperparameters. These include the number of LSTM layers $N_{\text{layer}}$, the embedding dimension $D_{\text{embed}}$, the hidden layer size $D_{\text{hidden}}$, and the number of iterations $N_{\text{iter}}$. Specifically, in the LSTM cell shown in Eq. 4, the weight matrices $\bm{W}_{f},\bm{W}_{i},\bm{W}_{o},\bm{W}_{c}\in\mathbb{R}^{D_{\text{embed}}\times(D_{\text{embed}}+D_{\text{hidden}})}$. A detailed summary of the parameters for the CNE neural network architecture and the simulation setup is provided in Table 2.